### v2

# Unification of the Gauge Theories

Abstract
Abstract
We treat the Christoffel coefficients as operators and introduce new mappings for quaternionic products to connect with the theory of electrodynamics in general spacetime. By utilizing the directional operator of the covariant derivative, we generalize the quaternionic mechanism to the theory of electrodynamics. We demonstrate that the Einstein equation allows selecting a constant term that aligns with the covariant derivative.
pacs:
03.67.Mn, 73.23.-b, 74.45.+c, 74.78.Na
Keywords: Quaternion, Electrodynamics, Christoffel Connections, Riemann’s Curvature Tensor, Gravity
I Introduction
After general relativity successfully explained the universe, many attempts were made to develop a geometric approach to electrodynamics. Unfortunately, most of these efforts were unsuccessful overduin ; ruiz ; wes ; cabral ; goenner ; lobo ; wess ; gross ; ponce ; williams ; wesso . Our current goal is to pursue this objective in a significantly different manner. We aim to incorporate the theory of gravity into the framework of electrodynamics. This integration may enhance our understanding of quantum gravity, a concept that physicists widely seek to realize. To validate the results of this work, we can analyze the components of Riemann’s curvature tensor. The pursuit of quantum gravity as a significant objective has led us to propose a modification in the dimensions of spacetime herbert ; kiefer ; gianluca ; gianlu ; gianluc ; carlo ; ashteka ; ashtekar . In addressing this challenge, we encounter local issues related to quaternionic production, particularly regarding the contraction of indices in tangent and cotangent spaces. However, the fundamental outcome remains consistent, regardless of the chosen space. These new perspectives will be instrumental in advancing our understanding of gravity theory. We examine two structures: the spacetime metric and the quaternionic metric, which involves quaternionic production under a new metric. As will be demonstrated, it is possible to transition from electrodynamics to gravity using quaternions. However, there is a fundamental difference between the two theories. Electrodynamics is a comprehensive theory consisting of two components, described by the strength tensor in four-dimensional space. When extending electrodynamics to a gravity theory using quaternions, we must consider four components in a four-dimensional spacetime. This leads to the idea of component separation. The alteration in the structure of these components results in equations of motion, ultimately culminating in Bianchi’s identity weber ; misner ; wald ; weinberg ; invarno ; malcolm . We will show that space (geometry) has a physical entity. Therefore, equations of motion will have to be determined, which go beyond mathematical relations to the properties of the geometry of space. When space takes on a physical meaning, it must obey the laws of physics, and ”space characters behavior must be affected by the equations of motion”. One of the consequences of this is that Einstein’s equation of gravity is just one of the determining equations of geometry.
The outcome of this article could be a modified gravitational theory, which holds the Einstein equation by formally including additional conditions gray ; parisi ; faisal ; sang . The additional conditions added to Einstein’s equation are viewed strictly from a purely theoretical perspective, which geometry alone has never achieved. We will derive Einstein’s equation with these additional conditions and demonstrate their necessity. The interpretation of Einstein’s historical theorem represents one of the transformative consequences of his theory. Understanding how gauge fields operate within the framework of gravity is a significant achievement misner ; wald ; invarno . Inducing the behavior of gauge fields in gravity is an achievement. The last theorem is made more complex by Riemann’s four-component curvature tensor. Certainly, the theory of gravity possesses a gauge-theoretic property analogous to gauge field theory. The directional covariant derivative operator is a key innovation in this work. We will discuss its effects in a subsection, but currently, we will utilize only its results.
Clarification- It is important to clarify definitions and designations: (1) The vector and dual spaces, represented by $\varepsilon$ and $\varepsilon^{\star}$, are ordered and denoted as $\varepsilon^{\star}\otimes\varepsilon$. Other arrangements are not permitted. The notation $\mathsf{u}^{\alpha\beta\cdots}_{\mu\nu\cdots}=u^{\alpha\beta\cdots}_{\mu\nu\cdots}\,(\hat{e}^{\mu}\hat{e}^{\nu}\cdots)\otimes(\hat{e}_{\alpha}\hat{e}_{\beta}\cdots)$ indicates a multicomponent member. The notation is revised as follows: the components $\mathsf{u}^{\alpha\beta}_{\mu\nu}$ will now be expressed as $u^{\alpha\beta}_{\mu\nu}$. The interaction of indices occurs only from a vector space to the dual space and vice versa; in any case, it is from a vector space into the dual space.
Establishing- "$\mathsf{ast}"$ is the new mapping introduced by defining its result as the only direction-independent action of the vector on its relevant dual space,
$\ast:\ \varepsilon\vec{\times}\varepsilon^{\star}\sim\mathrm{is\ deffined}$
Also, the definition of the mapping as an action is:
$\ast:\ \varepsilon^{\star}\reflectbox{$\vec{\reflectbox{$\times$}}$}\varepsilon\sim\mathrm{is\ deffined}$
In this way, $\varepsilon^{\star}\vec{\times}\varepsilon$ and $\varepsilon\reflectbox{$\vec{\reflectbox{$\times$}}$}\varepsilon^{\star}$ lead to invalid results. We can also represent this with $e_{\alpha}\vec{\times}e^{\beta}$ and so on. We introduce a Gamma operator that is directional:
$\hat{\Gamma}\ast\varepsilon^{\star}\rightarrow\varepsilon^{\star},\ \ \ \ \varepsilon\ast\hat{\Gamma}\rightarrow\varepsilon,$
Here, we will not mention the direction of the action. Therefore, operator $\hat{\Gamma}$ obtains a structure of $\varepsilon^{\star}\times\varepsilon$ and consequently $\hat{\Gamma}\propto\varepsilon^{\star}\times\varepsilon$. We point out that the operation of $\hat{\Gamma}$ is the contraction. In summary, $\hat{\Gamma}$ is an operator in the space of indices.
Due to assumption 1, the contraction of the indices occurs up to the allowed cases; $\Sigma_{\alpha}\ \hat{e}_{\alpha}\ast\ (\hat{e}^{\mu}\hat{e}^{\nu})=\Sigma_{\alpha}\ (\delta^{\mu}_{\alpha}\hat{e}^{\nu}+\hat{e}^{\mu}\delta^{\nu}_{\alpha})$. We assign a Lorentz index of $\hat{\Gamma}_{\mu}$ to the operator.
Metric- According to the definition of "$\mathsf{ast}"$ mapping, the metric is a mapping of the form;
$\mathsf{g}^{\mathrm{up}}:\varepsilon^{2}\times\varepsilon^{\star}\rightarrow\varepsilon$ (1)
and
$\mathsf{g}_{{}_{\mathrm{down}}}:\varepsilon^{\star}\leftarrow\varepsilon\times\varepsilon^{\star 2}$ (2)
The symbol "$\mathbf{g}"$ is the general metric element and comprises $\varepsilon\times\varepsilon$ or $\varepsilon^{\star}\times\varepsilon^{\star}$. Equations 1 and 2 show that the metric (with a repeated index) does not obey the rules of the operator, and it is clear that any other interpretation of equations 1 and 2 would be wrong for the action of the metric. It can be demonstrated that the metric effect does not influence the calculation of the components. Following the definition of "$\mathsf{ast}"$ mapping, the operator $\hat{\Gamma}_{\kappa}$ is represented as follows;
$\hat{\Gamma}_{\kappa}\equiv\Sigma_{\alpha\beta}(\Gamma^{\alpha}_{\beta\kappa}\ e^{\beta}e_{\alpha})^{\dagger}=\Gamma^{\alpha}_{\beta\kappa}\ e^{\alpha}e_{\beta}$ (3)
We have considered the inverted version of $\hat{\Gamma}_{\kappa}$ as the operator due to the compatibility of the Christoffel connections with the operator’s role. Furthermore, except for the "$\kappa"$ index in $\hat{\Gamma}_{\kappa}$, its other two indices are repeated and dummy. According to assumption 2, $\hat{\Gamma}_{\kappa}\equiv\hat{\Gamma}_{\kappa}\ e^{\kappa}$ (without summation) completes the missing part of relation equation 3. From this follows (a connection coefficient of the first kind):
$\Gamma^{\alpha}_{\beta\kappa}\mathsf{g}_{{}_{\mathrm{down}}}\rightarrow\Gamma^{\mu}_{\beta\kappa}\ g_{\mu\nu}\equiv\Gamma_{\beta\kappa\nu},$
the mapping is, of course, not an operator. Due to the specificity of contracting indices, we will remove the indicator from the top of the letters. With the notation of
$e_{\alpha}\rightarrow\ <\alpha|$
$e^{\alpha}\rightarrow\ |\alpha>$
The above statements and results become easier to understand with the new representation because the $\hat{\Gamma}_{\zeta}$ operator yields "$\hat{\Gamma}_{\kappa}=\Sigma_{\alpha,\beta}\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|\equiv\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|"$. The "$\mathsf{ast}"$ mapping becomes more formal when referring to the ”bra- ket” representations. Moreover, the representation corresponding to the ”bra-ket” extends our possibilities for future calculations. The $\alpha$ and $\beta$ are dummy indices in the $\Gamma^{\alpha}_{\beta\kappa}|\alpha><\beta|$-operator. They are used, among other things, for the gamma operator effect;
$\hat{\Gamma}_{\kappa}\ast\mathsf{u}_{\lambda}$ $=\Gamma^{\alpha}_{\beta\kappa}\ |\alpha><\beta|\lambda>u_{\lambda}=\Gamma^{\alpha}_{\beta\kappa}\ \delta_{\beta\lambda}\ u_{{}_{\mathrm{vacancy}}}|\alpha>$
$=\Gamma^{\alpha}_{\lambda\kappa}u_{\alpha},$ (4)
In this process, the box is mated with the alpha vector. That is, the box takes the alpha index. Hence,
$\mathsf{u}^{\lambda}\ast\hat{\Gamma}_{\kappa}=u^{\lambda}\Gamma^{\alpha}_{\beta\kappa}\ <\lambda|\alpha><\beta|=\Gamma^{\lambda}_{\beta\kappa}u^{\beta}.$ (5)
We emphasize that in changing the $\mathsf{u}_{\lambda}$ notation, the position of the index $\lambda$ in "$u_{\lambda}"$ is the main, but the name of the index changes in $\ast$-multiplication. However, we should use the "$\mathsf{ast}"$ mapping relations as simply as possible.
The relationships between the quaternion generators $\hat{q}_{1}=I$, $\hat{q}_{2}=J$, and $\hat{q}_{3}=K$ are defined within a Euclidean metric and adhere to the following properties: $\hat{q}_{i}\cdot\hat{q}_{j}=\delta_{ij}$ and $\hat{q}_{i}\times\hat{q}_{j}=\Sigma_{k}\epsilon_{ijk}\hat{q}_{k}$. If we include $\hat{q}_{0}=\Im$ (as the identity), we obtain a quaternion represented by the Lorentz quantities (parameters of the Lorentz index). These are coupled and are termed “$q$ vectors.” Lorentz quantities, which carry the Lorentz index, are represented using quaternions. We have two representations of the Lorentz index: the four-vector and the q-vector. We will use the notation $\mathtt{A},\ \mathtt{B},\cdots$ to denote the quaternions. Quaternions are typically constructed from four-vectors, denoted as $A_{\mu}=(A_{0},\mathbf{A})$. They can be expressed as $\mathtt{A}=iA_{0}\Im+A_{k}\hat{q}^{k}\equiv(\imath\mathtt{A}_{0},\mathtt{A})_{q}$, where the $q$-variable signifies the expression $(\imath\mathtt{A}_{0},\mathtt{A}_{k}\hat{q}^{k})$. Bold letters denote the spatial components of the four vectors. The four-vector framework is based on the spacetime metric, while the Euclidean metric pertains to the quaternion representation. The configuration significantly affects quaternionic products. We also analyze the Lorentz index, which connects physical parameters to the quaternion generator. We can generalize the range of coupled quantities from vectors to tensors. For the $q$-vectors $\hat{\Gamma}$ and $\mathtt{C}$, the extended Grassman quaternion multiplication """"$\triangleright$"" ▷ " (with $\ast$-product), is defined as follows andre ; walker ; kansu ; stefano
$\hat{\Gamma}\triangleright\mathtt{C}=(\hat{\Gamma}_{0}\ast\mathtt{C}_{0}-\hat{\Gamma}\ast\mathtt{C}\mathtt{C},\hat{\Gamma}\vee\mathtt{C}+\hat{\Gamma}\wedge\mathtt{C}\mathtt{C})_{q},$ (6)
where $\wedge$ and $\vee$ are the vector products under the "$\mathsf{ast}"$ mapping:
$\hat{\Gamma}\wedge\mathtt{C}\mathtt{C}=\epsilon^{ijk}(\hat{\Gamma}_{i}\ast\mathtt{C}_{j})\hat{q}_{k}.$
and
$(\hat{\Gamma}\vee\mathtt{C})_{k}=\hat{\Gamma}_{0}\ast\mathtt{C}_{k}-\hat{\Gamma}_{k}\ast\mathtt{C}_{0}.$
According to the above, $\vee$ and $\wedge$ add the spatial and temporal components to the above two commutators;
i) The commutator associated with the $\ast$-product is represented by the symbol $[\hat{\Gamma},\mathtt{C}]_{\ast}=\hat{\Gamma}\ast\mathtt{C}-\mathtt{C}\ast\hat{\Gamma}$.
ii) Moreover, we can introduce a new commutator, i.e. the commutator associated with the multiplication of the quaternion $[\mathtt{B},\mathtt{C}]_{\triangleright}=\mathtt{B}\triangleright\mathtt{C}-\mathtt{C}\triangleright\mathtt{B}$ containing the product """"$\ast$"" ∗ ", when at least one of them is operator. There are two distinct definitions of commutators.
Based on the operator rule for the connection coefficients $\hat{\Gamma}$, we introduce a new covariant derivative,
$\Delta_{\mu}=\partial_{\mu}\Im-e\hat{\Gamma}_{\mu}$ (7)
where "$e"$ is the coupling coefficient constant with a value of one. As mentioned earlier, $\hat{\Gamma}_{\mu}$ contains two coupling indices. Due to the operators $\hat{\Gamma}$’s and the $\ast$-product $\Delta_{\mu}$ is a semi-complete operator. Equation 7 with the $\ast$-product has the expected effects
$\Delta_{\mu}\ast\mathtt{B}^{\lambda}_{\kappa}$ $=\partial_{\mu}\mathtt{B}^{\lambda}_{\kappa}-e\Gamma^{\alpha}_{\mu\kappa}\mathtt{B}^{\lambda}_{\alpha}$
$=:\mathtt{B}^{\lambda}_{\kappa,\mu}$
The multiplications used in this paper refer to the $\ast$-product. Accordingly, the effect described above dictates the following action,
$[\Delta_{\mu},(\mathtt{B}\mathtt{C})]_{\ast}$ $=\mathtt{B}[\Delta_{\mu},\mathtt{C}]_{\ast}+[\Delta_{\mu},\mathtt{B}]_{\ast}\mathtt{C}$ (8)
It follows that we have
$[\Delta_{\mu},\mathtt{B}]_{\ast}$ $=\partial_{\mu}\mathtt{B}-e[\hat{\Gamma}_{\mu},\mathtt{B}]_{\ast}$
$\equiv\mathtt{B}_{;\mu}$ (9)
Quaternions- If we define $g^{\alpha\beta}\hat{\Gamma}_{\beta}$ directly, we can write
$g^{\alpha\beta}\Delta_{\beta}\ast\mathtt{B}_{\lambda}=g^{\alpha\beta}\mathtt{B}_{\lambda,\beta}=g^{\alpha\beta}\mathtt{B}_{\lambda;\beta}\equiv\mathtt{B}_{\lambda}^{\ ;\alpha}$
The covariant derivative is not complete. The incompleteness of the derivative operator is evident from equations I and 5.
$\mu$ and $\zeta$ are free indices in $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}$. This means that $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}$ ends up being a second rank tensor; In other words, $\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta}\propto\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}$ actually has the two free indices $\mu$ and $\zeta$ that can be considered as free indices; $\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}\equiv\Gamma^{\alpha}_{\mu\zeta}B_{\alpha}\ |\mu,\zeta>$. Equation 3, and our calculations in the ”bra-ket” notation show that,
$\hat{\Gamma}_{\nu}\ast\hat{\Gamma}_{\mu}$ $=\Gamma^{\alpha}_{\nu\beta}|\beta><\alpha|\Gamma^{\kappa}_{\mu\lambda}|\lambda><\kappa|+\Gamma^{\alpha}_{\nu\beta}|\alpha><\beta|\mu>$
×$\ \ \ \ \times\Gamma^{\kappa}_{\mu\lambda}|\lambda><\kappa|$
$=\Gamma^{\alpha}_{\nu\beta}\Gamma^{\kappa}_{\mu\lambda}|\beta>(\delta_{\alpha}^{\lambda})<\kappa|+\Gamma^{\alpha}_{\nu\beta}|\alpha>(\delta_{\beta\mu})$
×$\ \ \ \ \times\Gamma^{\kappa}_{\mu\lambda}|\lambda><\kappa|$
$=\Gamma^{\alpha}_{\nu\beta}\Gamma^{\kappa}_{\mu\alpha}|\beta><\kappa|+\Gamma^{\alpha}_{\nu\mu}\Gamma^{\kappa}_{\alpha\lambda}|\lambda><\kappa|,$
furthermore
$(\hat{\Gamma}_{\nu}\ast\hat{\Gamma}_{\mu})^{\dagger}$ $=(\Gamma^{\alpha}_{\nu\beta}\Gamma^{\kappa}_{\mu\alpha}+\Gamma^{\alpha}_{\nu\mu}\Gamma^{\kappa}_{\alpha\beta})|\beta><\kappa|)^{\dagger}$
$=(\Gamma^{\alpha}_{\nu\beta}\Gamma^{\kappa}_{\mu\alpha}+\Gamma^{\alpha}_{\nu\mu}\Gamma^{\kappa}_{\alpha\beta})|\kappa><\beta|,$
so the inverted version becomes
$(\hat{\Gamma}_{\nu}\ast\hat{\Gamma}_{\mu})\ast\mathtt{B}_{\zeta}=\Gamma^{\alpha}_{\nu\zeta}\Gamma^{\kappa}_{\mu\alpha}B_{\kappa}+\Gamma^{\alpha}_{\nu\mu}\Gamma^{\kappa}_{\alpha\zeta}B_{\kappa}.$
It also has the free indices, $\mu$, $\nu$ and $\zeta$. Also,
$\hat{\Gamma}_{\nu}\ast(\hat{\Gamma}_{\mu}\ast\mathtt{B}_{\zeta})$ $=\Gamma^{\alpha}_{\nu\beta}|\alpha><\beta|\ast(|\mu,\zeta>\Gamma^{\kappa}_{(\mu\zeta)}B_{\kappa})$
$=\Gamma^{\alpha}_{\nu\zeta}\Gamma^{\kappa}_{\mu\alpha}B_{\kappa}+\Gamma^{\alpha}_{\nu\mu}\Gamma^{\kappa}_{\alpha\zeta}B_{\kappa}.$ (10)
In addition,
$\hat{\Gamma}_{\kappa}\ast(\hat{\Gamma}_{\lambda}\ast\mathtt{B}_{\mu})=(\hat{\Gamma}_{\kappa}\ast\hat{\Gamma}_{\lambda})\ast\mathtt{B}_{\mu}.$
confirms that $\Delta_{\mu}$ is an incomplete operator, satisfying the condition:
$\Delta_{\mu}\ast(\mathtt{B}_{\lambda}\mathtt{C}_{\kappa})=(\Delta_{\mu}\ast\mathtt{B}_{\lambda})\mathtt{C}_{\kappa}+\mathtt{B}_{\lambda}(\Delta_{\mu}\ast\mathtt{C}_{k}).$
In particular,
$[\hat{\Gamma}_{\mu},\hat{\Gamma}_{\nu}]_{\ast}=(\Gamma^{\zeta}_{\mu\lambda}\Gamma^{\kappa}_{\nu\zeta}-\Gamma^{\zeta}_{\nu\lambda}\Gamma^{\kappa}_{\mu\zeta})(|\kappa><\lambda|).$
The free indices after transposition can be seen. In this way, $(\Gamma^{\zeta}_{\mu\lambda}\Gamma^{\kappa}_{\nu\zeta}-\Gamma^{\zeta}_{\nu\lambda}\Gamma^{\kappa}_{\mu\zeta})(|\kappa><\lambda|)$ as an operator contains the two free indices $\kappa$ and $\lambda$. It becomes a simplicial $[\hat{\Gamma}_{\mu},\hat{\Gamma}_{\nu}]^{\kappa}_{\ast\ \lambda}$ which has free $\lambda$ and $\kappa$ indices. Next, we follow the computation of $[\Delta_{\alpha},\Delta_{\beta}]_{\ast}$. We compute the case without torsion,
$\hat{\Gamma}_{\kappa;\alpha;\beta}$ $=\partial_{\beta}\partial_{\alpha}\hat{\Gamma}_{\kappa}-e\partial_{\beta}(\hat{\Gamma}_{\alpha}\ast\hat{\Gamma}_{\kappa})-e\Gamma^{\mu}_{\beta\alpha}\partial_{\mu}\hat{\Gamma}_{\kappa}$
$-e\hat{\Gamma}_{\beta}\partial_{\alpha}\ast\hat{\Gamma}_{\kappa}+e^{2}\hat{\Gamma}_{\beta}\ast(\hat{\Gamma}_{\alpha}\ast\hat{\Gamma}_{\kappa}),$
also
$\hat{\Gamma}_{\kappa;\beta;\alpha}$ $=\partial_{\alpha}\partial_{\beta}\hat{\Gamma}_{\kappa}-e\partial_{\alpha}(\hat{\Gamma}_{\beta}\ast\hat{\Gamma}_{\kappa})-e\Gamma^{\mu}_{\alpha\beta}\partial_{\mu}\hat{\Gamma}_{\kappa}$
$-e\hat{\Gamma}_{\alpha}\partial_{\beta}\ast\hat{\Gamma}_{\kappa}+e^{2}\hat{\Gamma}_{\alpha}\ast(\hat{\Gamma}_{\beta}\ast\hat{\Gamma}_{\kappa}),$
consequently
$\hat{\Gamma}_{\kappa;\alpha;\beta}-\hat{\Gamma}_{\kappa;\beta;\alpha}=e\hat{R}_{\alpha\beta}\ast\hat{\Gamma}_{\kappa}.$ (11)
Eq.11 means
$[\Delta_{\alpha},\Delta_{\beta}]_{\ast}=e\hat{R}_{\alpha\beta}.$ (12)
We are in a situation where the four components of Riemann’s curvature tensor have merged into a single operator with two distinct subspaces. Then, we will look at the roles of the components. With the action role and equation I, and for space with torsion, the new covariant derivative satisfies the following commutation relation, $[\Delta_{\alpha},\Delta_{\beta}]_{\ast}=e\hat{R}_{\alpha\beta}+e^{2}(\Gamma^{\lambda}_{\alpha\beta}-\Gamma^{\lambda}_{\beta\alpha})\hat{\Gamma}_{\lambda}$. But the effect of $\hat{R}_{\alpha\beta}$ on variable $\mathtt{B}_{\mu}$ is now,
$\hat{R}_{\alpha\beta}\ast\mathtt{B}_{\mu}$ $=(\partial_{\alpha}\hat{\Gamma}_{\beta}-\partial_{\beta}\hat{\Gamma}_{\alpha}-e\hat{\Gamma}_{\alpha}\ast\hat{\Gamma}_{\beta}+e\hat{\Gamma}_{\beta}\ast\hat{\Gamma}_{\alpha})\ast\mathtt{B}_{\mu}$
$=R^{\lambda}_{\ \mu\alpha\beta}B_{\lambda}|\mu,\alpha,\beta>.$ (13)
Another result is accessible; $[\Delta_{\mu},\mathsf{A}]_{\ast}$ reach to
$[\Delta_{\mu},A^{\alpha}_{\beta}]_{\ast}=\partial_{\mu}A^{\alpha}_{\beta}-g\Gamma^{\lambda}_{\beta\mu}A^{\alpha}_{\lambda}+g\Gamma^{\alpha}_{\lambda\mu}A^{\lambda}_{\beta},$
and
$[\Delta_{\mu},\mathsf{A}]^{\alpha}_{\ast\ \beta}$ $=(\partial_{\mu}\mathsf{A}-g\hat{\Gamma}_{\mu}\ast\mathsf{A}+g\mathsf{A}\ast\hat{\Gamma}_{\mu})^{\alpha}_{\beta}$
$=\partial_{\mu}A^{\alpha}_{\beta}-g\Gamma^{\kappa}_{\mu\zeta_{i}}A^{\cdots\zeta_{j}\cdots}_{..\zeta_{i-1}\kappa\zeta_{i+1}..}<\beta|\vec{\zeta}_{i}><\vec{\zeta}_{j}|\alpha>$
$-A_{\cdots\zeta_{i}\cdots}^{..\zeta_{j-1}\kappa\zeta_{j+1}..}\Gamma^{\zeta_{i}}_{\kappa\mu}<\beta|\vec{\zeta}_{i}><\vec{\zeta}_{j}|\alpha>.$
Now, considering relations I and I, we can write:
lemma- The covariance derivative effect is independent of entering the components.
Since $[\hat{\Gamma}_{\mu},\mathtt{A}_{\nu}]_{\ast}+[\mathtt{A}_{\mu},\hat{\Gamma}_{\nu}]_{\ast}=0$ , we take $\mathtt{A}$ to be as the gauge fields, then the change of the tensor of the field strength is
$\mathcal{E}_{\mu\nu}$ $=\partial_{\mu}\mathtt{A}_{\nu}-\partial_{\nu}\mathtt{A}_{\mu}$
$=\Delta_{\mu}\ast\mathtt{A}_{\nu}-\Delta_{\nu}\ast\mathtt{A}_{\mu}$ (14)
II Electrodynamics
In this section, we present a mechanism whereby the components of gauge fields $\mathbf{A}$ are coupled with the generators of the quaternions ($\mathbf{A}\rightarrow\mathtt{A}$). A connection between the components of the covariant derivative and the quaternion generators is similarly established. The quaternion variables obtained in this way help us to reconstruct the strength tensor of the fields. Due to the semi-completeness of the covariant derivative operator, we will introduce a new quaternion commutator to obtain the equation of motion and conclude the theory.
Based on the information provided, we will define new, commonly used variables and operators for quatrains. Starting from the tangent spaces, we introduce the operators and vectors as $q$-variables, $z_{t}=(\imath\mathtt{A}^{0},\mathtt{A}^{i}\hat{q}_{i})$ and $D_{t}=(\imath\Delta_{0},\mathbf{\Delta}_{i}\hat{q}^{i})$. The position of the indices is important here. $\mathtt{A}$ vectors are the fields; $\mathtt{A}\mathtt{A}$ is the vector potential and $\mathtt{A}_{0}=\phi$ represents the scalar case. There are two more i.e. $z_{c}=(\imath\mathtt{A}_{0},\mathtt{A}_{i}\hat{q}^{i})$ and $D_{c}=(\imath\Delta^{0},\mathbf{\Delta}^{i}\hat{q}_{i})$.
We show that Maxwell’s equations are accessible with augmented quaternions. The given content confirms the above definitions of the $q$-vectors. We emphasize that the quaternion mapping creates two separate spaces: the tangent and the cotangent $q$-vectors. The measures of two spaces are linked by the new matrix elements, rather than the Minkowski metric. Our calculations show that the law of direct and crossed multiplication between $q$-vectors is given by $H^{\mu\nu}=\left(\begin{array}[]{cccc}1&\imath&\imath&\imath\\ -\imath&-1&1&-1\\ -\imath&-1&-1&1\\ -\imath&1&-1&-1\\ \end{array}\right)$, where
$H^{\mu\nu}=\eta^{\mu\nu}+\left(\begin{array}[]{cccc}0&\imath&\imath&\imath\\ -\imath&0&1&-1\\ -\imath&-1&0&1\\ -\imath&1&-1&0\\ \end{array}\right):=\eta^{\mu\nu}+H^{\mu\nu}_{d},$
$\eta^{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$ is the Minkowski metric. With the help of the matrix $H^{\mu\nu}$ we define the inverse matrix
$\tilde{H}^{\mu\nu}=(H^{\mu\nu})^{T}=\left(\begin{array}[]{cccc}1&-\imath&-\imath&-\imath\\ \imath&-1&-1&1\\ \imath&1&-1&-1\\ \imath&-1&1&-1\\ \end{array}\right)=\eta^{\mu\nu}+\tilde{H}^{\mu\nu}_{d},$
We also introduce a new tensor, denoted as $n_{k}^{\ ij}$, derived from the behavior of quaternions. This tensor has the following properties: it is antisymmetric concerning the exchange of indices, and can be used to raise and lower indices; it is spatial, although it does not involve counting, and it includes a temporal index. The $n_{k}^{\ ij}$-symbol is used when its indices are all on the same row. When all indices are in a row, it resembles the Levi-Civita symbol. We will identify the location of the temporal index later.
Since the product of the quaternions is a quaternion, the new q-variable in the quaternion space can be named as $\mathsf{I}$, so we obtain $D_{t}\triangleright\ z_{c}=(\imath\mathsf{I}_{0},\mathsf{I}_{i}\hat{q}^{i})\equiv\mathsf{I}_{c}$. Considering the indices, we obtain
$\mathsf{I}_{c}=(\imath\mathsf{I}_{0},\mathsf{I})$
(using the upper indices), our calculations show that
$\mathsf{I}_{c}$ $=(\mathtt{A}_{\nu,\mu}\ \eta^{\mu\nu},\mathtt{A}_{\nu,\mu}\ H_{d}^{\mu\nu})$
based on this result and Eq.I, we arrive at
$\imath\mathsf{I}_{0}$ $=\mathtt{A}_{\nu,\mu}\ \eta^{\mu\nu}$ (15)
and
$\mathsf{I}_{k}$ $=\mathtt{A}_{\nu,\mu}\ H_{d}^{\mu\nu}$
are available with the following definitions:
$\mathsf{I}_{k}$ $=\imath\mathcal{E}_{0k}+\frac{1}{2}n^{\ \ ij}_{k}\mathcal{E}_{ij}$
$\equiv\imath\mathcal{E}_{0k}+D_{k},$
where $\mathcal{E}_{\mu\nu}$ is given by Eq.I. The equation of motion describes the action of the covariant derivative on the corresponding quaternion variables and vice versa. Due to the incompleteness of the derivative effects, to obtain the equation of motion, we must define a new commutator as follows: $[\ ,\ ]_{\triangleright}:\ \mathsf{Q}\times\mathsf{Q}\rightarrow\mathsf{Q}$ which is the same as $[\mathtt{B},\mathtt{C}]_{\triangleright}=(\eta^{\mu\nu}[\mathtt{B}_{\mu},\mathtt{C}_{\nu}]_{\ast},[\mathtt{B},\mathtt{C}]^{\vee}_{\ast}+[\mathtt{B}\mathtt{B},\mathtt{C}\mathtt{C}]^{\wedge}_{\ast})$. Due to the electrodynamics in Minkowski spacetime, the commutator of the motion has the following form (under the matrix $\tilde{H}$):
$[(\imath g^{0\zeta}\Delta_{\zeta},g^{s\zeta}\mathbf{\Delta}_{\zeta}),\mathsf{I}_{c}]^{\tilde{H}}_{\triangleright}$ $=(g^{\mu\zeta}\mathsf{I}_{\nu;\zeta}\eta^{\mu\nu},\mathsf{I}_{\nu;\mu}\ \tilde{H}_{d}^{\mu\nu})=\mathsf{J}_{c},$
$\mathsf{J}_{c}=(\imath\mathsf{J}_{0},\mathsf{J}_{i}\hat{q}^{i})$ is the quaternion form of current density. So the equation of motion expands to,
$g^{\mu\zeta}\mathsf{I}_{\nu;\zeta}\eta^{\mu\nu}$ $=\imath\mathsf{j}_{0},$
$\mathsf{I}_{\nu;\mu}\ \tilde{H}_{d}^{\mu\nu}$ $=\mathsf{j}_{k}$ (16)
[…]
$\Delta_{\nu}\ast\mathtt{B}_{\mu}=\mathtt{B}_{\mu,\nu}\equiv\mathtt{B}_{\mu;\nu}$ (17)
If $\mathtt{B}$ carries no upper indices, then $\Delta_{\nu}\ast\mathtt{B}_{;\mu}=\mathtt{B}_{;\mu;\nu}$ and $\Delta_{\mu}\ast g_{\kappa\lambda}=0$, $\Delta_{\mu}\ast g^{\kappa\lambda}\neq 0$. Therefore, equations 8 and II based on equation 17 lead to the following equation:
$g^{0\zeta}\mathsf{I}_{0\ ;\zeta}-\imath g^{s\zeta}\mathcal{E}_{0s;\zeta}-g^{s\zeta}D_{s;\zeta}=\imath\mathsf{J}_{0},$ (18)
$-\imath g^{0\zeta}(\imath\mathcal{E}_{0k;\zeta}+D_{k;\zeta})-n_{ks}^{\ \ \ l}g^{s\zeta}(\imath\mathcal{E}_{0l;\zeta}+D_{l;\zeta})$ (19)
$+\imath g^{k\zeta}\mathsf{I}_{0\ ;\zeta}=\mathsf{J}_{k},$ (20)
and finally, we arrived at
$-\imath g^{0\zeta}\mathsf{I}_{0\ ;\zeta}-g^{s\zeta}\mathcal{E}_{0s;\zeta}$ $=\mathsf{J}_{0},$
$g^{s\zeta}D_{s;\zeta}$ $=0,$
$\imath g^{k\zeta}\mathsf{I}_{0\ ;\zeta}+g^{0\zeta}\mathcal{E}_{0k;\zeta}-n_{ks}^{\ \ \ l}g^{s\zeta}D_{l;\zeta}$ $=\mathsf{J}_{k},$
$g^{0\zeta}D_{k;\zeta}+n_{ks}^{\ \ \ l}g^{s\zeta}\mathcal{E}_{0l;\zeta}$ $=0$ (21)
From the equations of motion, it follows that $g^{s\zeta}D_{s;\zeta}=0$, which gives a Bianchi identity (without time index): $\frac{1}{2}n_{s}^{\ ij}g^{s\zeta}\mathcal{E}_{ij;\zeta}\rightarrow\frac{1}{2}\varepsilon^{sij}\mathcal{E}_{ij;s}=0$. Moreover, the fourth equation is $g^{0\zeta}D_{k;\zeta}+n_{ks}^{\ \ \ l}g^{s\zeta}\mathcal{E}_{0l;\zeta}=\frac{1}{2}n_{k}^{\ ij}\mathcal{E}_{ij}^{\ \ ;0}+n_{k}^{\ \ sl}\mathcal{E}_{0l;s}=0$ which gives another Bianchi identity: $\frac{1}{2}n_{k0}^{\ \ \ ij}\mathcal{E}_{ij}^{\ \ \ ;0}+n_{k}^{\ s0l}\mathcal{E}_{0l;s}=\frac{1}{2}\varepsilon^{k0ij}\mathcal{E}_{ij;0}+\varepsilon^{ks0l}\mathcal{E}_{0l;s}=0$.
Since $n_{ks}^{\ \ \ l}g^{s\zeta}D_{l;\zeta}=\frac{1}{2}\varepsilon_{ksl}g^{s\zeta}\varepsilon^{lij}\mathcal{E}_{ij;\zeta}=g^{s\zeta}\mathcal{E}_{ks;\zeta}$, equation II is thus
$-g^{s\zeta}\mathcal{E}_{0s;\zeta}=\mathsf{J}_{0}+\imath g^{0\zeta}\mathsf{I}_{0\ ;\zeta},$
$g^{0\zeta}\mathcal{E}_{0k;\zeta}-g^{s\zeta}\mathcal{E}_{ks;\zeta}=\mathsf{J}_{k}-\imath g^{k\zeta}\mathsf{I}_{0\ ;\zeta}$
to summarize,
$g^{\mu\zeta}\mathcal{E}_{\mu\nu;\zeta}$ $=\mathtt{J}_{\nu},$ (22)
where
$\mathtt{J}_{\nu}=\mathsf{J}_{\nu}+\imath\eta_{\nu\mu}g^{\mu\zeta}\mathsf{I}_{0\ ;\zeta}.$ (23)
The above relations are simplified as follows:
$\mathsf{I}_{0}^{\ ;0}-\mathsf{I}_{s}^{\ ;s}=\imath\mathsf{j}_{0}$
$-\imath\mathsf{I}_{k}^{\ ;0}+\imath\mathsf{I}_{0}^{\ ;k}-n_{ki}^{\ \ j}\mathsf{I}_{j}^{\ ;i}=\mathsf{j}_{k}$
From Eqs.11 or 12, we get the relation $n_{ki}^{\ \ j}\mathsf{I}_{j}^{\ ;i;k}=\frac{e}{2}n_{ki}^{\ \ j}\hat{R}^{ik}\ast\mathsf{I}_{j}$ and $\imath(\mathsf{I}_{k}^{\ ;k;0}-\mathsf{I}_{k}^{\ ;0;k})=\imath e\hat{R}^{k0}\ast\mathsf{I}_{k}$. If we substitute these points, we obtain
$\mathsf{j}_{0}^{\ ;0}+\mathsf{j}_{k}^{\ ;k}$ $=-\imath\eta_{\mu\nu}g^{\mu\zeta}\mathsf{I}_{0;\zeta}^{\ \ ;\nu}+\imath e\hat{R}^{k0}\ast\mathsf{I}_{k}-\frac{e}{2}n_{ki}^{\ \ j}\hat{R}^{ik}\ast\mathsf{I}_{j}$
$=-\imath\eta_{\mu\nu}g^{\mu\zeta}\mathsf{I}_{0;\zeta}^{\ \ ;\nu}$
$+\frac{\imath e}{2}(n_{k}^{\ ij}\hat{R}^{k0}\ast\mathcal{E}_{ij}-n_{ki}^{\ \ j}\hat{R}^{ik}\ast\mathcal{E}_{0j})$
$-\frac{e}{4}(4\hat{R}^{k0}\ast\mathcal{E}_{0k}+n_{ki}^{\ \ j}n_{j}^{\ mn}\hat{R}^{ik}\ast\mathcal{E}_{mn})$
However, it can be seen that $n_{ki}^{\ \ j}n_{j}^{\ mn}\hat{R}^{ik}\ast\mathcal{E}_{mn}=\varepsilon_{kij}\varepsilon^{jmn}\hat{R}^{ik}\ast\mathcal{E}_{mn}=2\hat{R}^{nm}\ast\mathcal{E}_{mn}=2\hat{R}_{\nu\mu}\ast\mathcal{E}^{\mu\nu}-4\hat{R}_{0\mu}\ast\mathcal{E}^{\mu 0}$. Moreover, the second and third terms are zero because the equations of motion $n_{k}^{\ ij}\hat{R}^{0k}\ast\mathcal{E}_{ij}+n_{ki}^{\ \ j}\hat{R}^{ik}\ast\mathcal{E}_{0j}=0$. This result is independent of the physical conditions and has nothing to do with the sources. Substituting the results, we get
$\mathsf{j}_{0}^{\ ;0}+\mathsf{j}_{k}^{\ ;k}+\imath\eta_{\mu\nu}g^{\mu\zeta}\mathsf{I}_{0;\zeta}^{\ \ ;\nu}=-\frac{e}{2}\hat{R}^{\nu\mu}\ast\mathcal{E}_{\mu\nu}$ (24)
Now we derive equation 22 and from the fact that
$g^{\nu\kappa}\Delta_{\kappa}\ast g^{\mu\zeta}\mathcal{E}_{\mu\nu;\zeta}$ $=g^{\nu\kappa}\Delta_{\kappa}\ast\mathtt{J}_{\nu},$
we have: $g^{\mu\zeta}\mathcal{E}_{\mu\nu;\zeta}=g_{\nu\beta}\mathcal{E}^{\alpha\beta}_{\ \ ;\alpha}$ On the other hand, $g_{\nu\beta}\mathcal{E}^{\alpha\beta}_{\ \ ;\alpha}$ and the action of the derivative operator on the $\nu$ set is proportional to the low index set. From this follows: $g^{\nu\kappa}\Delta_{\kappa}\ast g_{\nu\beta}\mathcal{E}^{\alpha\beta}_{\ \ ;\alpha}=\mathcal{E}^{\alpha\beta}_{\ \ \ ;\alpha;\beta}$. The same logic applies to the $g^{\nu\kappa}\Delta_{\kappa}\ast\mathtt{J}_{\nu}=\mathtt{J}_{\nu;\kappa}g^{\kappa\nu}$. With these interpretations of equation I, we can write: ;$\mathcal{E}^{\mu\nu}_{\ \ ;\mu;\nu}=-\mathcal{E}^{\nu\mu}_{\ \ ;\nu;\mu}+e\hat{R}_{\mu\nu}\ast\mathcal{E}^{\mu\nu}$, and finally, we have
;$\mathcal{E}^{\mu\nu}_{\ \ ;\mu;\nu}=\frac{e}{2}\hat{R}_{\mu\nu}\ast\mathcal{E}^{\mu\nu}$ (25)
According to equations 24, 25, and the above, the continuity equation is as follows:
$\mathsf{j}_{0}^{\ ;0}+\mathsf{j}_{k}^{\ ;k}+\imath\eta_{\mu\nu}g^{\mu\zeta}\mathsf{I}_{0;\zeta}^{\ \ ;\nu}=\frac{e}{2}\hat{R}^{\mu\nu}\ast\mathcal{E}_{\mu\nu},$
But, $\mathcal{E}_{\mu\nu}^{\ \ ;\mu;\nu}=A_{\nu;\mu}^{\ \ ;\mu;\nu}-A_{\mu;\nu}^{\ \ ;\mu;\nu}=e\hat{R}^{\mu\nu}A_{\nu;\mu}$ and based on the relation in I, we reach $\mathcal{E}_{\mu\nu}^{\ \ ;\mu;\nu}\rightarrow-eR_{\mu}^{\ \lambda\mu\nu}A_{\lambda;\nu}-eR_{\nu}^{\ \lambda\mu\nu}A_{\mu;\lambda}=-eR_{\nu}^{\ \lambda\nu\mu}\mathcal{E}_{\mu\lambda}=0$, which means that the condition $\mathtt{J}_{\nu}^{\ ;\nu}=0$ as continuity equation is in agreement with reference lobo . Inserting equation I, equation 22 becomes
$g^{\mu\zeta}(\Delta_{\mu}\ast A_{\nu}-\Delta_{\nu}\ast A_{\mu})_{;\zeta}$ $=\mathtt{J}_{\nu},$
similar to the result of arbabb , or
$g^{\mu\zeta}\Delta_{\zeta}\ast\Delta_{\mu}\ast A_{\nu}-g^{\mu\zeta}\Delta_{\zeta}\ast\Delta_{\nu}\ast A_{\mu}$ $=\mathtt{J}_{\nu},$
but $g^{\mu\zeta}A_{\mu;\nu;\zeta}=g^{\mu\zeta}(\Delta_{\nu}\ast\Delta_{\zeta}+e\hat{R}_{\nu\zeta})\ast A_{\mu}$, which gives us
$g^{\mu\zeta}\Delta_{\zeta}\ast\Delta_{\mu}\ast A_{\nu}-eg^{\mu\zeta}\hat{R}_{\nu\zeta}\ast A_{\mu}$ $=\mathtt{J}_{\nu}+g^{\mu\zeta}\Delta_{\nu}\ast\Delta_{\zeta}\ast A_{\mu}.$
Since $g^{\mu\zeta}\Delta_{\nu}\ast\Delta_{\zeta}\ast A_{\mu}=g^{0\zeta}A_{0;\zeta;\nu}+g^{s\zeta}A_{s;\zeta;\nu}$ and the given condition that $A^{\kappa}_{\ ;\kappa}=0=g^{\kappa\lambda}A_{\lambda;\kappa}=g^{0\kappa}A_{0;\kappa}+g^{s\kappa}A_{s;\kappa}$, we can derive that $g^{s\zeta}A_{s;\zeta}=-g^{0\zeta}A_{0;\zeta}$. We obtain the following equation parker ; parke ; lindgren ; arbab ; cameron :
$g^{\mu\zeta}\Delta_{\zeta}\ast\Delta_{\mu}\ast A_{\nu}-eg^{\mu\zeta}R^{\alpha}_{\ \mu\nu\zeta}A_{\alpha}$ $=\mathtt{J}_{\nu},$
is, by substitution of the Ricci tensor, equivalent to
$g^{\mu\zeta}\Delta_{\zeta}\ast\Delta_{\mu}\ast A_{\nu}-e\mathrm{R}^{\ \ \alpha}_{\nu}A_{\alpha}$ $=\mathtt{J}_{\nu},$ (26)
which is
$g^{\mu\zeta}A_{\nu;\mu;\zeta}-e\mathrm{R}^{\ \ \alpha}_{\nu}A_{\alpha}$ $=\mathtt{J}_{\nu}.$
To reach the final goal (Eq.26), the relation $\mathtt{A}_{s;s}-\mathtt{A}_{0;0}=\mathtt{A}_{s}^{\ ;\zeta}g_{s\zeta}-\mathtt{A}_{0}^{\ ;\zeta}g_{0\zeta}$ must be established, and this is a precondition. With the above gauge, the continuity equation will be: $\mathsf{j}_{0}^{\ ;0}+\mathsf{j}_{k}^{\ ;k}=0$.
To determine the position of the time index in the new tensor $n_{k}^{\ ij}$, it is sufficient to define the following:
$G^{\mu\nu}=\frac{1}{2}\varepsilon^{\mu\nu\alpha\beta}\mathcal{E}_{\alpha\beta},$
so that $G^{\mu\nu}_{\ \ ;\nu}=0$. For the case $\mu=0$ the second equation of motion in Eq.II is now valid:
$G^{0k}_{\ \ ;k}=\frac{1}{2}\varepsilon^{0sij}\mathcal{E}_{ij;s},$ (27)
and for the case $\mu\neq 0$ we get the fourth equation of motion in Eq.II
$G^{l\zeta}_{\ \ ;\zeta}=\frac{1}{2}\varepsilon^{l0ij}\mathcal{E}_{ij;0}+\varepsilon^{li0j}\mathcal{E}_{0j;i}=0$ (28)
which gives the position of the time index in the comparison.
III Conclusions
We demonstrate that proving gravitational issues, such as dark matter, Einstein’s equation can’t take any additional terms. Modifying Einstein’s equation is possible only by vanishing covariant derivative terms.
IV Acknowledgments
The author would like to thank Shahrekord University for providing the research grant.
V Geometry
In this section, the gauge fields are replaced by the Christoffel connections $\mathsf{A}$ to $\Gamma$. This change does not affect the definition of the covariant derivative operator. Using the same method of the previous section and multiplying the four components of the Riemann’s curvature tensor into two two-component subspaces, we define the quaternion variables as $z_{c}=(\imath\hat{\Gamma}_{0},\hat{\Gamma}_{i}\hat{q}^{i})$ and $D_{c}=(\imath\Delta^{0},\mathbf{\Delta}^{i}\hat{q}_{i})$. It is clear that $\hat{\Gamma}^{0}=g^{0\zeta}\hat{\Gamma}_{\zeta}$, and this is a general case. If we set the above points and emphasize the position of the indicators, we can write
$\mathsf{I}_{c}$ $=(\hat{\Gamma}_{\nu,\mu}\eta^{\mu\nu},\hat{\Gamma}_{\nu,\mu}\tilde{H}_{d}^{\mu\nu})$
With the help of these preliminary remarks, and similar to the relation in 15, we arrive at the definitions
$\imath\mathsf{I}_{0}$ $=\hat{\Gamma}_{\nu,\mu}\eta^{\mu\nu}$
and the equation
$\mathsf{I}_{k}=\hat{\Gamma}_{\nu,\mu}\tilde{H}_{d}^{\mu\nu},$
is presented in shorthand form as follows:
$\mathsf{I}_{k}$ $=\imath\mathcal{F}_{0k}+\frac{1}{2}n_{k}^{\ ij}\mathcal{F}_{ij}$
$=\imath\mathcal{F}_{0k}+\mathfrak{D}_{k}$
The new notation $\mathcal{F}_{\mu\nu}$ (Eq.I) serves to visually distinguish this variable from that of the previous section, $\mathcal{E}_{\mu\nu}$. As a result, all previous relationships and their arrangements remain intact with this transformation. Thus, the relation in 18 will be the equation of motion for the new fields $\mathcal{F}_{\mu\nu}$. By entering the quaternion current density and separating the main equations into real and imaginary components, we have
$-\imath g^{0\zeta}\mathsf{I}_{0\ ;\zeta}-g^{s\zeta}\mathcal{F}_{0s;\zeta}$ $=\mathsf{J}_{0},$
$g^{s\zeta}D_{s;\zeta}$ $=0,$
$\imath g^{k\zeta}\mathsf{I}_{0\ ;\zeta}+g^{0\zeta}\mathcal{F}_{0k;\zeta}-n_{ks}^{\ \ \ l}g^{s\zeta}D_{l;\zeta}$ $=\mathsf{J}_{k},$
$g^{0\zeta}D_{k;\zeta}+n_{ks}^{\ \ \ l}g^{s\zeta}\mathcal{F}_{0l;\zeta}$ $=0$ (29)
The connection between the geometric tensor dynamics and sources is clear. Comparing the results from the previous section, it is evident that the properties of the gauge field strength tensor apply to both components of Riemann’s curvature tensor. These performances emphasize that the Christoffel connections have the gauge property. Equations 27 and V give the source less equations:
$\frac{1}{2}\varepsilon^{0kij}\hat{R}_{ij;k}=0.$
and from equation 28
$\frac{1}{2}\varepsilon^{l0ij}\hat{R}_{ij;0}+\varepsilon^{li0j}\hat{R}_{0j;i}=0$ (30)
Eq.30 represents the Bianchi identities and can be derived from $\hat{G}^{\mu\nu}=\frac{1}{2}\varepsilon^{\mu\nu\alpha\beta}\hat{R}_{\alpha\beta}$. For the Bianchi identities, the sub-components of the released space behave independently. But, in the second relation, which is independent of the source, the behavior of the second Bianichi identity is intertwined with the basis of the subcomponents. The remaining two relations of the basic equation of motion (Eq.V) are as follows:
$g^{\mu\zeta}\hat{R}_{\mu\nu;\zeta}$ $=\mathtt{J}_{\nu},$ (31)
which $\mathtt{J}_{\nu}$ in Eq.31, still gives the current density as Eq.23. together with $\hat{R}_{\mu\nu}^{\ \ ;\mu;\nu}=0$ and the condition $\mathtt{J}_{\nu}^{\ ;\nu}=0$. Besides the gauge condition, the precondition
$\eta_{\nu\mu}g^{\mu\zeta}\mathsf{I}_{0\ ;\zeta}=0,$
is also conceivable.
A new tensor, $G^{\kappa\lambda}_{\ \ \ \mu\nu}$, can be introduced here:
$G^{\kappa\lambda}_{\ \ \ \mu\nu}=\frac{1}{4}\varepsilon^{\kappa\lambda\alpha\beta}\varepsilon_{\mu\nu\theta\pi}R^{\theta\pi}_{\ \ \ \alpha\beta},$
which gets its special form:
$G^{\kappa\lambda}_{\ \ \ k\nu}$
$=-\frac{1}{2}\delta^{\lambda}_{\nu}R_{\alpha\beta}^{\ \ \ \alpha\beta}+R_{\nu\alpha}^{\ \ \ \lambda\alpha}.$
On the other hand, if we use the equation of motion (Eq.V)
$G^{\kappa\lambda}_{\ \ \ \mu\nu;\kappa}$ $=\frac{1}{4}\varepsilon^{\kappa\lambda\theta\pi}\varepsilon_{\mu\nu\alpha\beta}R^{\alpha\beta}_{\ \ \ \theta\pi;\kappa}=0$
$=\frac{1}{2}\varepsilon^{\alpha\lambda\theta\pi}\varepsilon_{\mu\nu\alpha\beta}\mathtt{J}^{\beta}_{\ \ \theta\pi}-\frac{1}{2}\delta_{\nu}^{\lambda}\ \mathcal{R}_{;\mu}$
$+R^{\lambda\alpha}_{\ \ \ \nu\alpha;\mu},$ (32)
$\mathcal{R}$ is the Ricci scalar. As far as we know
$\varepsilon^{\alpha\lambda\theta\pi}\varepsilon_{\mu\nu\alpha\beta}\mathtt{J}^{\beta}_{\theta\pi}$ $=-\delta^{\alpha}_{\alpha}\det{\left(\begin{array}[]{ccc}\delta^{\lambda}_{\mu}&\delta^{\lambda}_{\nu}&\delta^{\lambda}_{\beta}\\ \delta^{\theta}_{\mu}&\delta^{\theta}_{\nu}&\delta^{\theta}_{\beta}\\ \delta^{\pi}_{\mu}&\delta^{\pi}_{\nu}&\delta^{\pi}_{\beta}\end{array}\right)}\mathtt{J}^{\beta}_{\theta\pi}$
$=(4-1)!(-2\delta^{\lambda}_{\mu}\mathtt{J}^{\beta}_{\nu\beta}+2\delta^{\lambda}_{\nu}\mathtt{J}^{\beta}_{\mu\beta}-2\delta^{\lambda}_{\beta}\mathtt{J}^{\beta}_{\mu\nu})$
$:=\mathtt{S}^{\lambda}_{\mu\nu},$
so Eq.V becomes
$-\frac{1}{2}\delta^{\lambda}_{\nu}\ \mathcal{R}_{;\mu}+R^{\lambda\alpha}_{\ \ \ \nu\alpha;\mu}+\mathtt{S}^{\lambda}_{\mu\nu}=0,$ (33)
if one now sets $\mathtt{S}^{\lambda}_{\mu\nu}:=\mathtt{T}^{\lambda}_{\nu;\mu}$ and that based on the equation I, $-\frac{1}{2}\delta^{\lambda}_{\nu}\ \mathcal{R}_{;\mu}+R^{\lambda\alpha}_{\ \ \ \nu\alpha;\mu}=[\Delta_{\mu},\frac{1}{2}\delta^{\lambda}_{\nu}\ \mathcal{R}+R^{\lambda\alpha}_{\ \ \ \nu\alpha}]_{\ast}$, one finds a more complete form of 33,
$-\frac{1}{2}\delta^{\lambda}_{\nu}\ \mathcal{R}+R^{\lambda\alpha}_{\ \ \ \nu\alpha}=-\mathtt{T}^{\lambda}_{\nu}+\mathtt{C}^{\lambda}_{\nu}$
certainly
$\mathtt{C}^{\lambda}_{\nu;\mu}=0$ (34)
is like the famous case $\mathtt{C}_{\lambda\nu}=-\Lambda g_{\lambda\nu}$. It turns out that $\mathtt{C}_{\lambda\nu}=f_{\lambda\nu}(\mathbf{g})$ (for example $(e^{\kappa\mathbf{g}})_{\lambda\nu}$, $\mathbf{g}$ is the matrix for the metric). $\mathtt{C}_{\lambda\nu}$, is a geometrical quantity, not matter. It can be considered general background matter, which can be absorbed by the tensor $\mathtt{T}_{\lambda\nu}$ and become zero.
In the case of lower indices,
$-\frac{1}{2}g^{\lambda\beta}g_{\beta\nu}\ \mathcal{R}+g^{\lambda\beta}R^{\ \alpha}_{\beta\ \nu\alpha}=-g^{\lambda\beta}\mathtt{T}_{\beta\nu}+g^{\lambda\beta}\mathtt{C}_{\beta\nu}$ (35)
or
$\mathrm{R}_{\lambda\nu}-\frac{1}{2}g_{\lambda\nu}\ \mathcal{R}+\mathtt{C}_{\lambda\nu}=\mathtt{T}_{\lambda\nu}$ (36)
In the final equation, the characters of geometry and matter are distinguished, from one another. The fact that in Einstein’s theory of gravitation, Einstein’s term (in agreement with the covariant derivative) is not the only way to choose $\mathtt{C}_{\lambda\nu}$ is another result of this mechanism. We must also consider constants that indicate the limitations of spacetime to solve the equation of gravity. It is a question of finding the constant from equation 36 because it is not a unique quantity. Changing any principal constant in equation 36 (agreement with covariant derivative) leads us into a new universe. Among the possible options for the constant $\mathtt{C}_{\lambda\nu}$, equation 36 makes a perturbation theorem. What can be obtained, however, is the non-uniqueness of Einstein’s famous theorem.
VI Discussion
Here, we have used a quaternion mechanism to form new vectors. We have provided $q$-vectors as physical quantities—such as events in the manifold of spacetime—that correspond to the position of the Lorentz indices and are coupled to the generators of quaternion algebra. By introducing the operator $\hat{\Gamma}_{\mu}$ and the mapping "$\mathsf{ast}"$ ($\ast$-product), we then succeeded in obtaining the theory of gauge fields in a general spacetime. We introduce the new antisymmetric tensor $n_{k}^{\ ij}$ and use the equations of motion of gauge theory to obtain the Bianchi identities. We have also shown that the Bianchi identities are physical relations, not mathematical structures. As we proceeded, we replaced the gauge fields with the Christoffel coefficients and followed the improved quaternion mechanism to access the theory of gravity. Our calculations indicate that with this substitution and additional degrees of freedom for Riemann’s curvature tensor, the theory of gravitation remains valid, and the Einstein constant term is essential for the equations of motion. The results of gravity are much higher than those of electrodynamic theory, and the reason for this difference is Riemann’s curvature four-component tensor. As for the Bianchi identities in gravity, we have concluded that the identities are part of the equations of motion. One of the most significant outcomes of deriving the equation of gravity is the formalization of the constant term, referred to as Einstein’s constant term, in the Einstein equation. We have also established that the constant term is exclusively geometric. Having decomposed the four-component Riemann’s curvature tensor into one with two component subsections, another result of this work is to emphasize the independent behavior of the two-component subsections of the curvature tensor. The result of the present work is a commentary on parisi ; faisal and many other references and matches the physical data with the solutions of the Einstein equation, with many corrections to the Einstein theory of gravity using the unprincipled modification approach to the Einstein equation (modifications that do not apply to equation 34). By analyzing the components of Riemann’s curvature tensor, we can develop a new approach to quantum gravity. The conclusion of this article is that any manipulation of Einstein’s equation is incorrect and must comply with the condition ”the covariant derivative equals zero.”